Please help me determine if this is linear or not.
Determine whether the differential equation is linear . y'+ycosx=x^9y^9
Thanks.

Answer by Edwin McCravy(20056)

A linear differential equation is of the form

y' + P(x)*y = Q(x)

y'+ycosx=x^9y^9

No it's not linear because the right side is not a 
function of x only.

The differential equation 

y'+ycosx=x^9

is linear but not

y'+ycosx=x^9y^9

Edwin